MalayaJanae:

The map of a walking trail is drawn on a coordinate grid with three points of interest. The trail starts at R(−1, 4) and goes to S(5, 4) and continues to T(5, −2). The total length of the walking trail ____ is units

XxAngelGirlxX:

RS only changes x coordinates so x= 3-(-2)=? ST only changes y coordinates so y= 4-(-1)=? add x and y: x+y = ?